Find y if the slope between (-9,y) and (-18,14) is -1
Answer by stanbon(75887)

Find y if the slope between (-9,y) and (-18,14) is -1
-----------------------
(y-14)/(-9--18) = -1
----
(y-14)/9 = -1
-----
y-14 = -9
y = 5
----------
Cheers,
Stan H.
---------